OpenStudy (anonymous):

I really need some help! Can someone help me factor this? x^6-y^6

OpenStudy (hoblos):

x^6-y^6 = (x^3-y^3)(x^3+y^3) = (x-y)(x^2 +xy+y^2)(x+y)(x^2 -xy +y^2)

OpenStudy (anonymous):

Nice work.